OpenStudy (anonymous):

Find fog(2) when f(x)=-x^2 + 4X and g(x) = 2 ^x -1

OpenStudy (anonymous):

Put g(x) for ALL x's in f(x)

OpenStudy (anonymous):

thanks ! can you help me get a numerical answer i got 21 and I'm unsure if it is correct

OpenStudy (anonymous):

sure, can you confirm the that g(x)= 2^x-1

OpenStudy (anonymous):

You get 3 but then when i plug in for f(x) i get 21 but someone else told me it was three so i am unsure if i am right or if they are right

OpenStudy (anonymous):

Just type the question again

OpenStudy (anonymous):

Find fog(2) when f(x)=-x^2 + 4X and g(x) = 2 ^x -1, im trying to figure out if the final answer is 21

OpenStudy (anonymous):

\[is g(x)=2^{x} - 1\]

OpenStudy (anonymous):

what you have typed is pretty confusing

OpenStudy (anonymous):

yes it is that formula you wrote but it is g(x)

OpenStudy (anonymous):

\(g(2)=2^2-1=3\) find \(f(3)\)

OpenStudy (anonymous):

thanks, i got that part now i need to know if the final answer is 21

OpenStudy (anonymous):

since \(f(x)=-x^2 + 4x\) just do f(3) which means you plug in 3 for all the x's like so:\[f(3)=-(3)^2+4(3)~~~~\implies~~~~f(3)=-9+12~~~~\implies~~~~f(3)=3\] i hope it makes sense now! :) @wonderstruck

OpenStudy (anonymous):

yayyyy ! im so glad! lol thanks so much !! :DDD

OpenStudy (anonymous):

youre welcome! :D